1. Let f(n)2  = f(n +1) be a recurrence relation. Given f(0) = 2, solve.

2. Let  1+ \sum_{i=0}^{n-1}f(i) = f(n+1) be a recurrence relation. Given f(0) = 1, f(1) = 1 and n 1, solve.
Answer #1

1. = 8 2 . = (f(n))? f(n+1). f(0) = 2 f(1) = f(0+1) = (f(0) = 22 f(2) = flit) (F() (23) ? = 24 f(3) f(2+0) (f(2))? (24)2 = $f(+) = f(6+1) = 1 + f(0+ f(+ f(2) + f(3) + f(4) + f(5) f16 + f(5) - 13 + 8 = 2 ) Thus f(8) = f(7) + f(6) = 24 + 13 = 34 Look